# Thinker: Learning to Think Fast and Slow

### v1

@@ Abstract
Recent studies show that the reasoning capabilities of Large Language Models (LLMs) can be improved by applying Reinforcement Learning (RL) to question-answering (QA) tasks in areas such as math and coding. With a long context length, LLMs may learn to perform search, as indicated by the self-correction behavior observed in DeepSeek R1. However, this search behavior is often imprecise and lacks confidence, resulting in long, redundant responses and highlighting deficiencies in intuition and verification. Inspired by the Dual Process Theory in psychology, we introduce a simple modification to the QA task that includes four stages: Fast Thinking, where the LLM must answer within a strict token budget; Verification, where the model evaluates its initial response; Slow Thinking, where it refines the initial response with more deliberation; and Summarization, where it distills the refinement from the previous stage into precise steps. Our proposed task improves average accuracy from 24.9% to 27.9% for Qwen2.5-1.5B, and from 45.9% to 49.8% for DeepSeek-R1-Qwen-1.5B. Notably, for Qwen2.5-1.5B, the Fast Thinking mode alone achieves 26.8% accuracy using fewer than 1000 tokens, demonstrating substantial inference efficiency gains. These findings suggest that intuition and deliberative reasoning are distinct, complementary systems benefiting from targeted training.
@@ 1 Introduction
    Verification: The agent evaluates the correctness of the initial answer using a large token budget. If verified, it is accepted as the final answer.
@@
Experimental results validate our approach: relative to the QA task, the Thinker task yields consistent gains across diverse math benchmarks, with average relative performance gains of 11.9% for Qwen2.5-1.5B models and 8.50% for DeepSeek-R1-Distill-Qwen-1.5B models. Furthermore, our analysis reveals a notable reduction in reflection patterns, suggesting more direct reasoning. In summary, the proposed Thinker task offers the following key strengths:
@@ 2 Background
where $a\sim\pi_{\theta}(\cdot\mid x)$, and $R$ is the reward function, such as a binary function that returns $1$ if the extracted answer from $a$ matches the ground-truth answer $y^{*}$, and 00 otherwise.
@@ 3.1 Task Description
Step 2 - Verification. In the second step, the agent is prompted to verify whether the fast answer $y_{\text{fast}}$ is correct, and must output either Yes or No. The response in this step is restricted to a relatively long maximum generation length (e.g., 6000 tokens). The reward $R_{\text{verify}}$ in this step is defined as a weighted binary function based on the correctness of the verification. Specifically, let $y_{\text{verify}}$ denote the extracted answer; then:
$R_{\text{verify}}=\begin{cases}(1-p_{\text{fast-acc}})\cdot 1\{y_{\text{verify}}=\texttt{Yes}\}&\text{if $y_{\text{fast}}=y^{*}$,}\\ p_{\text{fast-acc}}\cdot 1\{y_{\text{verify}}=\texttt{No}\}&\text{otherwise},\\ \end{cases}$ (3)
@@
where $x_{\text{fast}}$ is the prompt in Fast Thinking step (i.e., the initial prompt), $\log P(a_{\text{summary}}\mid x_{\text{fast}})$ is the log probability of the summary response under the Fast Thinking prompt, and $c$ is a scalar hyperparameter. In experiments, we found that the agent sometimes still degenerates to give a very short answer despite the log probability term. To mitigate this, we gate the reward to 00 if the length of the generated response is less than a low threshold (e.g., 300 tokens).
@@ 3.2 Training with the Thinker Task
Training LLMs with the Thinker task requires particular considerations regarding reward propagation. Since the reward at each step is specific to that step alone, it should not be propagated backward to earlier steps. This implies that the discount factor between steps should be set to 00, while that within each step should be high (e.g., $1$) to enable effective credit assignment over tokens. The Thinker task defines only the RL environment and imposes no restrictions on the choice of algorithm or model architecture, allowing compatibility with any standard RL method (e.g., PPO) and LLM.
@@ 4 Related Work
RL and Reasoning in LLMs. A large number of studies have demonstrated the effectiveness of applying RL to enhance the reasoning capabilities of LLMs guo2025deepseek ; zeng2025simplerl ; hu2025open ; deepscaler2025 . Our work builds on these efforts by decomposing the QA task into the four-step Thinker task. It has been observed that LLMs trained with RL can produce inefficient CoT chen2024not , leading to overthinking sui2025stop . Our work relates to strategies for controlling generation length (often termed token budgets). Examples include dynamic token budgets that scale with problem complexity han2024token , or user-defined budgets aggarwal2025l1 ; xu2025scalable . For instance, Muennighoff et al. muennighoff2025s1 utilize token budget controls during the generation of CoT data for supervised fine-tuning (SFT). While these approaches primarily focus on token budget control within a single response generation or for SFT data preparation, our method introduces a structured, multi-step RL task explicitly designed to train distinct agent abilities. In addition, our work is related to methods for encouraging self-correction in LLMs kamoi2024can . For instance, methods like Self-Refine madaan2023self and Reflexion shinn2023reflexion primarily use prompt engineering and few-shot examples to enable agents to incorporate internal or external feedback for refining subsequent responses. However, unlike our work, these methods typically do not involve RL fine-tuning of the agent for these self-correction behaviors; instead, the correction capability is elicited at inference time through prompting.
@@ 5.1 Experimental Setup
To evaluate the Thinker task, we fine-tune two publicly available models: Qwen2.5-1.5B (Q1.5B) yang2024qwen2 and DeepSeek-R1-Distill-Qwen-1.5B (R1.5B) guo2025deepseek . While sharing a base architecture, R1.5B has undergone additional distillation using reasoning data from DeepSeek-R1, endowing it with stronger initial reasoning and search behaviors compared to the base Q1.5B. Training both models allows us to investigate the Thinker task’s impact on a model with foundational capabilities (Q1.5B) and its ability to further enhance a model already specialized for reasoning (R1.5B).
We fine-tune these models using RL on both the Thinker task and a standard QA task (serving as our baseline). For all experiments, we employ PPO schulman2017proximal . Key hyperparameters include a discount rate $\gamma=1$, GAE lambda $\lambda=1$, and a sampling temperature of $1$. No KL penalty against a reference policy was applied. The Fast Thinking and Summarization stages use a 1000-token budget, while Verification and Slow Thinking use a 6000-token budget. Most other hyperparameters and training details mirror those in Open-Reasoner-Zero, with details provided in Appendix A.
For the training data, we utilized the 129K math question-answering dataset provided by Open-Reasoner-Zero. Each training run for both the Thinker task and the baseline required approximately 7 days on two compute nodes, each equipped with 8 A100 GPUs. Our implementation, adapted from the Open-Reasoner-Zero codebase hu2025open , is publicly available at https://github.com/DualityRL/thinker-task.
@@ 5.2 Training Dynamics and Evaluation Results
The training performance is shown in Figure 3, measured by accuracy on questions in the training data. For the Thinker task, we plot the fast accuracy (the accuracy of the agent during the Fast Thinking step) and the final accuracy (the accuracy of the final answer). Note that this final accuracy is not directly comparable to that of the baseline, as the Thinker-task agent effectively has two attempts during training due to the transition function from Verification to Slow Thinking. Nonetheless, we observe that both its fast and final accuracy improve steadily, while the accuracy of the baseline plateaus rather quickly. This suggests that the Thinker-task agent is developing stronger underlying intuition for the task, potentially more so than just acquiring a generic search algorithm, which may contribute to its sustained improvement. We also observe that the fast accuracy exceeds the baseline’s accuracy in the Q1.5B model, indicating the effectiveness of the Fast Thinking mode.
@@
Figure 3: Accuracy on training set.
@@
Figure 4: Evaluation performance across seven common benchmarks.
The performance of the final models, evaluated on the aforementioned benchmarks and the additional CollegeMath benchmark, is detailed in Table 1. Due to the large number of questions in CollegeMath, we did not evaluate training checkpoints on this benchmark, only the final models. In addition to the Fast Accuracy and Final Accuracy of the Thinker-task agent and the accuracy of the fine-tuned baseline agent, Table 1 also includes results for: (1) the pretrained model (the model before any fine-tuning); (2) reported figures from Open-Reasoner-Zero hu2025open (labelled ‘ORZ’); and (3) reported figures from SimpleRL zeng2025simplerl . Both ORZ and SimpleRL utilize PPO to train agents on standard QA tasks.
From Table 1, we observe that the Thinker-task agent consistently performs better than the other methods across almost all benchmarks. For the Q1.5B model, the Thinker-Fast agent’s performance already surpasses that of the baseline, while the full Thinker-task agent achieves a 11.9% relative performance gain on average compared to this baseline. For the R1.5B model, the Thinker-Fast agent performs slightly worse than the baseline but still significantly outperforms the pretrained model. This result for Thinker-Fast is notable, suggesting a substantial efficiency gain, as it only requires 1000 token budget compared to the 8000 token budget of the pretrained model. The full Thinker-task model with R1.5B surpasses the baseline by an average relative increase of 8.50%. These results collectively suggest the benefits of decomposing the single-turn QA task into the proposed four-step Thinker task.
Table 1: Performance comparison across various mathematical reasoning benchmarks. Average (Avg.) scores are presented. All scores are Pass@1 accuracy (%) averaged over 16 samples. Top score in each benchmark column is bolded. Standard errors are provided in Appendix B.
@@
Baseline 57.98 3.33 3.33 21.46 24.54 34.38 17.78 36.21 24.88
Thinker 64.25 6.25 2.50 23.74 28.11 40.62 19.03 38.33 27.85
Thinker-Fast 61.60 6.25 2.50 26.39 24.78 35.94 18.66 37.85 26.75
@@
Thinker 87.02 35.62 27.71 36.08 54.21 81.72 33.23 42.77 49.80
Thinker-Fast 77.21 11.46 11.46 30.08 40.39 59.22 29.23 41.37 37.55
@@ 5.3 Analysis and Case Study
Reflection and Response Length. Beyond overall performance, we investigated whether the Thinker task encourages more direct reasoning with less overt self-reflection or self-doubt. We compiled a vocabulary list of terms commonly associated with self-reflection in Deepseek R1’s outputs (e.g., "wait," "however," "alternatively") and measured the average occurrence of these reflection patterns during training, with results shown in Figure 5b. We observed that the Thinker-task agent tends to use fewer such reflection patterns in its reasoning trace, suggesting it may be learning to reason more directly.
Paradoxically, despite reduced reflection patterns, the Thinker agent’s responses were generally longer than the baseline’s (Figure 5b). This increased length primarily stems from repeated self-verifications within the Verification step (e.g., "Let me check… Yes. Let me check again… Yes."), explaining the growing verification length in Figure 5b. This behavior is likely incentivized by our answer extraction focusing only on the last boxed output, encouraging reiteration and refinement if tokens remain. Future work could mitigate this by, for instance, permitting only a single boxed answer in the Verification step.
@@
Role of Summarization. Finally, we investigated the importance of the Summarization step through an ablation analysis. We trained a Q1.5B model using a modified Thinker task where the Summarization step was removed. The average reflection pattern count and response length for the Thinker task, both with and without the Summarization step, are presented in Figure 6. Without the Summarization step, the model exhibited a more frequent increase in reflection patterns, accompanied by highly fluctuating response lengths during training. This suggests that the Summarization step may contribute to stabilizing the learning process, potentially by discouraging degeneration into excessive reflection.
Removing the Summarization step negatively impacted Fast Thinking: average Fast Accuracy on benchmarks dropped from 26.75% (with summarization) to 24.84% (without). This suggests summarization, by potentially distilling concise reasoning, enhances intuition for Fast Thinking. Interestingly, final accuracy remained comparable (27.85% with vs. 27.84% without summarization), indicating that the Slow Thinking step could often compensate for the reduced fast accuracy. Detailed ablation results can be found in Appendix B.
@@
6 Limitation
This study has several limitations. Firstly, our experiments were confined to two 1.5B parameter models, and we did not investigate the scalability of the Thinker task to larger models due to computational constraints. Secondly, hyperparameters, such as token limits for each task stage, were set heuristically without systematic tuning, suggesting our reported results may not be optimal. Lastly, agents trained with the Thinker task generally produce slightly longer responses than standard QA agents, potentially impacting computational efficiency at inference if all stages are utilized.

### v2

@@ Abstract
Recent studies show that the reasoning capabilities of Large Language Models (LLMs) can be improved by applying Reinforcement Learning (RL) to question-answering (QA) tasks in areas such as math and coding. With a long context length, LLMs may learn to perform search, as indicated by the self-correction behavior observed in DeepSeek R1. However, this search behavior is often imprecise and lacks confidence, resulting in long, redundant responses and highlighting deficiencies in intuition and verification. Inspired by the Dual Process Theory in psychology, we introduce a simple modification to the QA task that includes four stages: Fast Thinking, where the LLM must answer within a strict token budget; Verification, where the model evaluates its initial response; Slow Thinking, where it refines the initial response with more deliberation; and Summarization, where it distills the refinement from the previous stage into precise steps. Our proposed task improves average accuracy from 25.6% to 27.3% for Qwen2.5-1.5B, and from 45.9% to 51.0% for DeepSeek-R1-Qwen-1.5B. Notably, for Qwen2.5-1.5B, the Fast Thinking mode alone achieves 25.2% accuracy using fewer than 1000 tokens, demonstrating substantial inference efficiency gains. These findings suggest that intuition and deliberative reasoning are distinct, complementary systems benefiting from targeted training. Additionally, we have open-sourced both the trained models and the source code.
@@ 1 Introduction
    Verification: The agent evaluates the correctness of the initial answer using a small token budget. If verified, it is accepted as the final answer.
@@
Experimental results validate our approach: relative to the QA task, the Thinker task yields consistent gains across diverse math benchmarks, with average relative performance gains of 6.7% for Qwen2.5-1.5B models and 11.1% for DeepSeek-R1-Distill-Qwen-1.5B models. Furthermore, our analysis reveals a notable reduction in reflection patterns, suggesting more direct reasoning. In summary, the proposed Thinker task offers the following key strengths:
@@ 2 Background
where $a\sim\pi_{\theta}(\cdot\mid x)$, and $R$ is the reward function, such as a binary function that returns $1$ if the extracted answer from $a$ matches the ground-truth answer $y^{*}$, and 0 otherwise.
@@ 3.1 Task Description
Step 2 - Verification. In the second step, the agent is prompted to verify whether the fast answer $y_{\text{fast}}$ is correct, and must output either Yes or No. The response in this step is restricted to a relatively short maximum generation length (e.g., 2000 tokens). The reward $R_{\text{verify}}$ in this step is defined as a weighted binary function based on the correctness of the verification. Specifically, let $y_{\text{verify}}$ denote the extracted answer; then:
[…]
@@
where $x_{\text{fast}}$ is the prompt in Fast Thinking step (i.e., the initial prompt), $\log P(a_{\text{summary}}\mid x_{\text{fast}})$ is the log probability of the summary response under the Fast Thinking prompt, and $c$ is a scalar hyperparameter. In experiments, we found that the agent sometimes still degenerates to give a very short answer despite the log probability term. To mitigate this, we gate the reward to 0 if the length of the generated response is less than a low threshold (e.g., 300 tokens).
@@ 3.2 Training with the Thinker Task
Training LLMs with the Thinker task requires particular considerations regarding reward propagation. Since the reward at each step is specific to that step alone, it should not be propagated backward to earlier steps. This implies that the discount factor between steps should be set to 0, while that within each step should be high (e.g., $1$) to enable effective credit assignment over tokens. The Thinker task defines only the RL environment and imposes no restrictions on the choice of algorithm or model architecture, allowing compatibility with any standard RL method (e.g., PPO) and LLM.
@@ 4 Related Work
RL and Reasoning in LLMs. A large number of studies have demonstrated the effectiveness of applying RL to enhance the reasoning capabilities of LLMs guo2025deepseek ; zeng2025simplerl ; hu2025open ; deepscaler2025 . Our work builds on these efforts by decomposing the QA task into the four-step Thinker task. It has been observed that LLMs trained with RL can produce inefficient CoT chen2024not , leading to overthinking sui2025stop . Our work relates to strategies for controlling generation length (often termed token budgets). Examples include dynamic token budgets that scale with problem complexity han2024token , or user-defined budgets aggarwal2025l1 ; xu2025scalable . For instance, Muennighoff et al. muennighoff2025s1 utilize token budget controls during the generation of CoT data for supervised fine-tuning (SFT). Concurrent works such as ThinkPrune hou2025thinkprune , Concise Reasoning fatemi2025concise , SR-Flow dong2025enhancing , and AdaptThink zhang2025adaptthink modify the reward or RL algorithm to encourage more efficient reasoning. While these approaches primarily focus on token budget control within a single response generation, our method introduces a structured, multi-step RL task explicitly designed to train distinct agent abilities independent of the underlying RL algorithm. In addition, our work is related to methods for encouraging self-correction in LLMs kamoi2024can . For instance, methods like Self-Refine madaan2023self and Reflexion shinn2023reflexion primarily use prompt engineering and few-shot examples to enable agents to incorporate internal or external feedback for refining subsequent responses. However, unlike our work, these methods typically do not involve RL fine-tuning of the agent for these self-correction behaviors; instead, the correction capability is elicited at inference time through prompting.
@@ 5.1 Experimental Setup
To evaluate the Thinker task, we fine-tune two publicly available models: Qwen2.5-1.5B (Q1.5B) yang2024qwen2 and DeepSeek-R1-Distill-Qwen-1.5B (R1.5B) guo2025deepseek . While sharing a base architecture, R1.5B has undergone additional distillation using reasoning data from DeepSeek-R1, endowing it with stronger initial reasoning and search behaviors compared to the base Q1.5B. Training both models allows us to investigate the Thinker task’s impact on a model with foundational capabilities (Q1.5B) and its ability to further enhance a model already specialized for reasoning (R1.5B). For the Q1.5B runs, we use three independent seeds and report the averaged results.
We fine-tune these models using RL on both the Thinker task and a standard QA task (serving as our baseline). For all experiments, we employ PPO schulman2017proximal . Key hyperparameters include a discount rate $\gamma=1$, GAE lambda $\lambda=1$, and a sampling temperature of $1$. No KL penalty against a reference policy was applied. The Fast Thinking and Summarization stages use a 1000-token budget, Verification uses a 2000-token budget, and Slow Thinking uses a 6000-token budget. Most other hyperparameters and training details mirror those in Open-Reasoner-Zero, with details provided in Appendix A.
For the training data, we utilized the 129K math question-answering dataset provided by Open-Reasoner-Zero. Each training run for both the Thinker task and the baseline required approximately 7 days on two compute nodes, each equipped with 8 A100 GPUs. Our implementation, adapted from the Open-Reasoner-Zero codebase hu2025open , is publicly available at https://github.com/stephen-chung-mh/thinker-task, which also includes the trained models.
@@ 5.2 Training Dynamics and Evaluation Results
The training performance is shown in Figure 3, measured by accuracy on the training data. For the Thinker task, we plot fast accuracy (from the Fast Thinking stage) and final accuracy (from the Slow Thinking stage). Note that the final accuracy is not directly comparable to the baseline’s, as the Thinker agent effectively has two attempts during training (though only one during testing). We observe that the Thinker agent’s fast and final accuracy improve steadily, while the baseline’s accuracy plateaus rather quickly. This may suggest that the Thinker agent is learning the underlying intuition of the task, which requires sustained training, as opposed to merely acquiring a generic search algorithm. We also observe that the fast accuracy of the Q1.5B model surpasses the baseline’s. Notably, while both metrics reflect a single attempt, the Fast Thinking mode achieves superior performance with a significantly smaller token budget (Figure 4), underscoring its efficiency.
Figure 4 illustrates the average response length during training, plotting both the Fast Thinking response length and the cumulative length across all four stages of the Thinker task. The two models exhibit distinct behaviors. For the Q1.5B model, which lacks strong inherent self-correction capabilities, the Thinker task’s cumulative length is greater than the baseline’s, as its structured stages introduce verification and refinement steps. Conversely, for the R1.5B model, which already possesses self-correction abilities, the Thinker agent’s total response length eventually becomes shorter than the baseline’s. This suggests that rather than simply adding overhead, the Thinker task refines the R1.5B model’s existing search process, pruning inefficiencies to achieve more token-efficient reasoning.
@@
Figure 3: Accuracy on training set. For the Qwen2.5-1.5B model (a), the shaded region represents the standard deviation across three independent seeds.
@@
Figure 4: Average response length on training set. For the Qwen2.5-1.5B model (a), the shaded region represents the standard deviation across three independent seeds.
(a) Qwen2.5-1.5B
(b) DeepSeek-R1-Distill-Qwen-1.5B
Figure 5: Evaluation performance across seven common benchmarks. For the Qwen2.5-1.5B model (a), the shaded region represents the standard deviation across three independent seeds.
Table 1: Performance comparison across various mathematical reasoning benchmarks. Average (Avg.) scores are presented. All scores are Pass@1 accuracy (%) averaged over 16 samples. Top score in each benchmark column is bolded. Standard errors and more statistical analysis are provided in Appendix B. Results for the Q1.5B Thinker and Baseline are averaged across three independent seeds.
@@
Baseline 59.82 4.10 2.43 20.52 26.05 35.36 19.25 37.42 25.62
Thinker 64.45 6.25 2.22 19.21 28.21 39.06 20.38 38.82 27.33
Thinker-Fast 59.82 4.58 1.25 21.28 24.52 34.53 17.85 37.58 25.18
@@
Thinker 88.51 38.96 26.67 37.41 55.49 83.59 34.77 42.46 50.98
Thinker-Fast 81.35 18.33 14.58 28.85 45.68 66.41 31.39 41.74 41.05
The performance of the final models is detailed in Table 1. These models correspond to the checkpoints, saved at 50-step intervals, that achieved the highest accuracy on a validation dataset. We evaluated these final models on the aforementioned benchmarks and the additional CollegeMath benchmark; due to its large size, CollegeMath was reserved for this final evaluation and was not used to assess intermediate checkpoints. For a comprehensive comparison, the table includes the Fast and Final Accuracy of our Thinker agent, the accuracy of the fine-tuned Baseline, and results for the original Pretrained model. Since evaluation is performed in inference mode, the Thinker agent does not require external verification, making its Final Accuracy directly comparable to the single-attempt baselines. Furthermore, we include reported figures from Open-Reasoner-Zero (ORZ) hu2025open and SimpleRL zeng2025simplerl , both of which also utilize PPO to train agents on standard QA tasks.
From Table 1, we observe that the Thinker agent consistently performs better than the other methods across almost all benchmarks. For the Q1.5B model, the Thinker-Fast agent’s performance is already close to that of the baseline, while the full Thinker agent achieves a 6.7% relative performance gain on average compared to this baseline. For the R1.5B model, the Thinker-Fast agent performs slightly worse than the baseline but still significantly outperforms the pretrained model. This result for Thinker-Fast is notable, suggesting a substantial efficiency gain, as it only requires 1000 token budget compared to the 8000 token budget of the pretrained model. The full Thinker-task model with R1.5B surpasses the baseline by an average relative increase of 11.1%. These results collectively suggest the benefits of decomposing the single-turn QA task into the proposed four-step Thinker task. To confirm these benefits extend to larger models, additional experiments on 7B models were conducted, yielding similar results as detailed in Appendix D.
To evaluate token efficiency, we analyze the Thinker agent on three benchmarks used in concurrent work on this topic (Table 2). The results show that our agent achieves superior performance across all three benchmarks with a competitive token count. The Thinker-Fast mode is also effective on easier datasets like MATH500 and AMC23, using substantially fewer tokens. This highlights a key advantage for deployment: the flexibility to select the appropriate mode based on the task’s complexity.
Table 2: Comparison with concurrent works fine-tuning R1.5B models on token efficiency benchmarks. Results from concurrent works are extracted from the respective papers.
Method MATH500 AIME24 AMC23
Acc. (%) Length Acc. (%) Length Acc. (%) Length
ThinkPrune hou2025thinkprune 83.2 1938 27.1 5631 73.2 3039
Concise Reasoning fatemi2025concise 81.0 1965 30.0 6752 69.4 2936
SR-FLOW dong2025enhancing 85.3 - 36.7 - 77.8 -
AdaptThink zhang2025adaptthink 82.0 1782 31.0 6679 - -
Baseline 86.2 2780 35.4 5778 72.8 3938
Thinker 88.5 2501 39.0 5597 83.6 3517
Thinker-Fast 80.9 600 18.1 853 66.9 751
@@ 5.3 Analysis and Case Study
Reflection and Response Length. Beyond overall performance, we investigated whether the Thinker task encourages more direct reasoning with less overt self-reflection or self-doubt. We compiled a vocabulary list of terms commonly associated with self-reflection in Deepseek R1’s outputs (e.g., "wait," "however," "alternatively") and measured the average occurrence of these reflection patterns during training, with results shown in Figure 6a. We observed that the Thinker agent tends to use fewer such reflection patterns in its reasoning trace, suggesting it may be learning to reason more directly.
A detailed breakdown of response length during training is presented in Figure 6b. The trends for the two agents diverge: after stabilizing, the baseline’s response length steadily increases, likely due to an increase in learned self-reflection. Conversely, the Thinker agent’s total response length gradually decreases. This improved efficiency is not due to a reduction in the length of individual stages, which remain consistent. Rather, it stems directly from the model’s rising Fast Thinking accuracy. As the agent’s intuition improves, it solves more problems correctly in the Fast Thinking stage, thereby allowing it to bypass the token-intensive Slow Thinking stage and significantly reduce its average token length.
@@
Role of Summarization. [^1] This ablation study used a 6000-token Verification step. The length was later reduced to 2000 tokens for the main experiments, as this did not significantly affect performance. Finally, we investigated the importance of the Summarization step through an ablation analysis. We trained a Q1.5B model using a modified Thinker task where the Summarization step was removed. The average reflection pattern count and response length for the Thinker task, both with and without the Summarization step, are presented in Figure 7. Without the Summarization step, the model exhibited a more frequent increase in reflection patterns, accompanied by highly fluctuating response lengths. This indicates that the Summarization step may contribute to stabilizing the learning process, potentially by discouraging degeneration into excessive reflection, though we cannot rule out that the observed divergence is due to stochastic variation.
Removing the Summarization step negatively impacted Fast Thinking: average Fast Accuracy dropped from 26.75% to 24.84%. This suggests summarization, by potentially distilling concise reasoning, enhances intuition for Fast Thinking. Interestingly, final accuracy remained comparable (27.85% with vs. 27.84% without summarization), indicating that the Slow Thinking step could often compensate for the reduced fast accuracy. Detailed ablation results can be found in Appendix B.